a SVD for A and A2 if 2. Perform 1 3 2 -1 0 0 2 _ A 4 -2 1 -4 0 2
Answer #1

SVD means singular value decomposition.

SVD for the given Matrix A is as follows.

Here we have to find eigenvalues and eigenvectors.

Solution: 2 1 -1 3 0 0 -1 2 A = 1 -2 4 1 -4 2 +4.A 15 7 -1 12 7 5 -2 -1 -2 -13 22 12 8 -13 21 Find Eigen vector for A A A A A

(i-6313 +966 2-26521 1681) = 0 (24-6313 +9662-26521+ 1681 Roots can be found using newton raphson method +Newton Raphson meth

+2. Eigenvectors for A 18.73127615 = 6.14651506 2.62859506 7.46582039 1 +3. Eigenvectors for A 2.35397647 -1.09159899 0.34944

-1.67201863 0.3328001 0.26160317 1 So, normalizing gives v4 = ( - 0.8386512, 0.16692589, 0.13121493, 0.50158006) 1.99369969 10.15538378 0.47625711 -0.21377373 -0.8386512 0.22869316 -0.20155155 -0.93766314 0.16692589 . V -0.86871823 0.39204305 -0.2727

Singular value matrix for the A^2.

2 1 -1 3 -1 2 0 2 1 -1 3 6 4 -19 13 -1 2 0 -2 0 0 1 2 -12 3 4 1 1 -2 4 7 -7 13 -25 3 1 1 -4 2 0 -4 2 1 0 1

Now to find the required singular value decomposition for A^2.

Is given below.

In this step, we are finding the eigenvalues and eigenvectors.

Solution: 2 1 -1 3 0 0 -1 2 A = 1 -2 4 1 -4 2 +4.A 15 7 -1 12 7 5 -2 -1 -2 -13 22 12 8 -13 21 Find Eigen vector for A A A A A

Roots can be found using newton raphson method +Newton Raphson method for x -1747x 3 + 368269x2- 2080803x+2825761 0 . x - 2.2

4. Eigenvectors for 2 2.22482285 -0.20567786 -1.20154187 V 0.55428461 1 = 0.827978442 +0.468992292+(-048023937)2+12 1.4615512

1st Solution 1502.87790279 0 0 0 38.76696922 0 238.3516052 0 0 0 15.43864 0 0 ,Σ= V3.54566916 0 1.88299473 0 C 0 0 1.49158401

Thanks,

Keep smiling and do CHEEGING.

Please do rate and review.